sove for x.
4x(x-1)-5x(x)=3

Answer by ewatrrr(24785)


Hi
solving for x
4x(x-1)-5x(x)=3
4x^2 - 4x -5x^2 = 3
x^2 +4x + 3 = 0
factoring
(x+3)(x+1) = 0 Note:SUM of the inner product(3x) and the outer product(1*x) =4x
(x+3)= 0
x = -3
(x+1) = 0
x = -1